OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).??????????????

OpenStudy (anonymous):

First, you need to get 4x-3y=-12 into slope-intercept form, which is y=mx+b. Can you do this?

OpenStudy (ahsome):

@pookie0799 , I am kinda busy so I cannot go step by step. I will later. Till then, what you need to do is find the \(m\) value of that equation, which is the slope Simply turn the equation into \(y=mx+b\) and it should be easy from there Then input that value, as well as the point onto this formula, which will give you your equation. \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

I know i got y=4x-8 ......but now what?

OpenStudy (ahsome):

Your \(m\) is 4. Since your equation is in the form \(y=mx+b\), your \(m\) value is \(4\)

OpenStudy (ahsome):

Now, sub that value, as well as your point, into \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

or you could sub (x,y) points into y=mx+b to solve for b..

OpenStudy (ahsome):

You could. But if you sub the value into the equation, then expand the bracketss, it will also give you the \(b\) value